# HAPNet: Toward Superior RGB-Thermal Scene Parsing via Hybrid, Asymmetric, and Progressive Heterogeneous Feature Fusion

### v1

@@ I-B Existing Challenges and Motivation
Additionally, while symmetric duplex encoders [20, 21, 10] are prevalently used, they may not be the optimal architectures for heterogeneous feature extraction. This limitation arises from the substantial inherent modality difference [22] between the RGB and thermal images, as well as the unique advantages provided by vision Transformers (ViTs) and convolutional neural networks (CNNs). RGB images, rich in global semantic cues [23], are well-suited for learning through ViTs, while both RGB and thermal images provide gradient-related details (local semantics) [9, 24], which can be more effectively learned through CNNs. Therefore, developing a novel asymmetric duplex encoder, with ViT and CNN employed separately in each branch, for RGB-T scene parsing, stands as an underexplored area of research deserving greater attention.
@@ II-A Single-Modal Scene Parsing
The advent of the fully convolutional network (FCN) [26] catalyzed a wave of research on CNN-based scene parsing networks, which typically utilize convolutions for pixel-wise classification and explore various modules and architectures to enhance overall performance. For instance, the DeepLab series [27, 28] leverages atrous spatial pyramid pooling to capture multi-scale context information, while the feature pyramid network (FPN) [29] effectively fuses multi-scale features for robust dense prediction. Nonetheless, CNNs suffer from inherent limitations, e.g., constrained receptive fields, which impede their effectiveness in comprehensively modeling global context.
Transformers [30], originally proposed for natural language processing, have revolutionized computer vision due to their capability to model long-range dependencies. As a representative example, segmentation Transformer (SETR) [31] introduces a ViT-based encoder for scene parsing, enabling long-range dependency modeling via self-attention mechanisms. Another approach, SegFormer [32], employs a hierarchical encoder to achieve multi-scale feature encoding within the Transformer framework, outperforming SETR in handling objects of varying sizes. Additionally, drawing upon earlier works [33, 34], the MaskFormer [35, 36] series formulates scene parsing as a mask classification problem. These networks utilize a Transformer decoder to refine object queries and generate pixel-level predictions. Given the impressive performance demonstrated by the MaskFormer architecture in dense prediction tasks, we adopt a similar network design for scene parsing in this study.
@@ II-B Data-Fusion Scene Parsing
To address this drawback, recent works [21, 9, 8, 10, 6, 20, 7] have resorted to more advanced, learnable feature fusion strategies. For instance, FEANet [21] and GMNet [9] introduce attention-based modules to dynamically select important heterogeneous features. Their subsequent research endeavor, CMX [8], employs average and max pooling operations along with the attention modules to recalibrate heterogeneous features at the spatial and channel dimensions. Similarly, CAINet [10] employs different attention modules to model global context and aggregate local semantics across various spatial scales of RGB-T feature maps. Moreover, the ABMDRNet series [6, 20] introduces novel modality-wise bridging-then-fusing frameworks, capable of reconstructing pseudo images of one modality using the features extracted from the other modality. This approach successfully reduces modality discrepancy and generates more discriminative fused features. Additionally, inspired by the masked auto-encoder (MAE) [38], CRM-RGBTSeg [7] randomly corrupts fixed-size regions and subsequently recovers the corrupted regions from the other modality, preventing over-reliance on either modality. In this article, we employ a novel asymmetric duplex encoder to capture both local spatial priors and global context from the RGB-T data. These heterogeneous features are progressively fused for more robust scene parsing.
@@ III-A Overall Workflow
Unlike previous works that utilize symmetric duplex encoder structures, we innovatively introduce an asymmetric, hybrid architecture (based on both a Transformer-based VFM and a lightweight CNN). It effectively combines the complementary strengths of both modalities, thereby achieving deeper cross-modal feature fusion and further improving the overall performance. This is realized by dividing the ViT encoder into four blocks and inserting the PHFI to form four encoding stages. The RGB image $\boldsymbol{I}^{R}\in\mathbb{R}^{H\times W\times 3}$ and its corresponding thermal image $\boldsymbol{I}^{T}\in\mathbb{R}^{H\times W\times 3}$ are first fed into the CSPD, extracting the cross-modal spatial prior $\boldsymbol{F}^{P}_{1}\in\mathbb{R}^{(\sum_{i=2}^{4}{\frac{HW}{{S_{i}}^{2}})}\times D}$, which serves as one of the inputs to HAPNet, where $S_{i}=2^{i+1}$ ($i\in[2,4]\cap\mathbb{Z}$) denotes the corresponding stride number. Afterwards, $\boldsymbol{I}^{R}$ is tokenized to form the context feature $\boldsymbol{F}^{V}_{1}\in\mathbb{R}^{\frac{HW}{16^{2}}\times D}$, which serves as the other input to HAPNet. $\boldsymbol{F}^{V}_{1}$ undergoes global context encoding in the following ViT blocks. $\boldsymbol{F}^{V}_{1}$ and $\boldsymbol{F}^{P}_{1}$ are then fused in a dual-path, progressive manner using the two key components of our developed PHFI: (1) the global-local context aggregator (GLCA) and (2) the complementary context generator (CCG). This design enables the RGB features to retain fine-grained local semantics while effectively capturing global context through the powerful VFM. Eventually, the fused features $\mathcal{F}^{F}=\{\boldsymbol{F}^{F}_{\frac{1}{S_{j}}}\in\mathbb{R}^{\frac{H}{S_{j}}\times\frac{W}{S_{j}}\times D}\}$, where $S_{j}=2^{i+1}$ ($j\in[1,4]\cap\mathbb{Z}$) from the four encoding stages are obtained and then fed into the subsequent mask classification decoder to generate semantic predictions $\boldsymbol{M}^{P}\in\mathbb{R}^{H\times W}$, which are also utilized by an auxiliary local semantics enhancement task during model training.
@@ III-B3 Progressive Heterogeneous Feature Integrator
where $\mathrm{MHA}(\cdot)$ represents the multi-head attention (MHA) operation (implemented based on multi-scale deformable attention [36] to reduce computations) and $\mathrm{LN}(\cdot)$ represents the layer normalization operation. Following common practices in attention modules such as [45, 43], we introduce a learnable coefficient $\kappa_{i}$ to dynamically adjust the weight of the MHA, enabling flexible fusion of the local semantics-enhanced features into the VFM. After the $i$-th ViT block, the output feature maps $\boldsymbol{F}^{V}_{i+1}$ with enriched global context are yielded. $\boldsymbol{F}^{V}_{i+1}$ is then fed into the CCG to fuse its global context with the spatial prior $\boldsymbol{F}^{P}_{i}$. In this process, $\boldsymbol{F}^{P}_{i}$ serves as the query, while $\boldsymbol{F}^{V}_{i+1}$ acts as the keys and values to perform MHA within the CCG as follows:
@@ III-E Loss Function
We use a combination of the pixel-wise cross-entropy (CE) loss $\mathcal{L}_{\text{cls}}$ for class prediction, along with the binary CE loss $\mathcal{L}_{\text{bce}}$ and the dice loss $\mathcal{L}_{\text{dice}}$ for mask prediction. For our local semantics enhancement task, we directly calculate the standard CE loss $\mathcal{L}_{\text{ce}}$ over its semantic predictions. Therefore, the overall loss function can be expressed as follows:
@@
Following [36], we set $\lambda_{\text{bce}}=5.0$, $\lambda_{\text{dice}}=5.0$, $\lambda_{\text{cls}}=2.0$, and $\lambda_{\text{ce}}=0.4$, respectively. Additionally, we employ bipartite matching [47] to determine the least-cost assignment during model training.
@@ IV-C Implementation Details
Our model was trained for 200 epochs on an NVIDIA RTX 3090 GPU, utilizing the AdamW optimizer [60] with an initial learning rate of $2\times 10^{-4}$ and a weight decay of $5\times 10^{-2}$. Following the common practice employed in [16], we apply a layer-wise learning rate decay of $0.9$ for our VFM encoder. For the experiments conducted on the NYU dataset, we employ RGB images and the HHA encoding of depth images as model input. We conduct all ablation studies on the widely utilized MFNet dataset, with the auxiliary task omitted to ensure training efficiency. The performance improvement (an increase in mIoU by $0.6\%$) between Sect. IV-D and Sect. IV-E demonstrate the effectiveness of the auxiliary task.
@@ IV-D Comparisons with SoTA Scene Parsing Networks
Specifically, our proposed HAPNet achieves the highest mIoU across all datasets, outperforming SoTA methods by $0.1$-$8.3\%$ on the MFNet dataset, $1.0$-$32.0\%$ on the PST900 dataset, and $2.4$-$28.9\%$ on the KP day-night dataset, respectively. These results demonstrate the robustness and effectiveness of HAPNet for RGB-T scene parsing in various scenarios, ranging from complex urban driving scenes to challenging underground scenes. Notably, HAPNet outperforms CRM-RGBTSeg [7], another SoTA Transformer-based approach based on the mask classification paradigm, demonstrating the efficacy of our adapter-based hybrid, asymmetric architecture in heterogeneous feature fusion. The qualitative results also demonstrate the superior capability of HAPNet in handling various scenes with challenging lighting conditions.

### v2

@@ I-B Existing Challenges and Motivation
Additionally, while symmetric duplex encoders [20, 21, 10] are prevalently used, they may not be the optimal architectures for heterogeneous feature extraction. This limitation arises from the substantial inherent modality differences [22] between the RGB and thermal images, as well as the unique advantages provided by vision Transformers (ViTs) and convolutional neural networks (CNNs). RGB images, rich in global semantic cues [23], are well-suited for learning through ViTs, while both RGB and thermal images provide gradient-related details (local semantics) [9, 24], which can be more effectively learned through CNNs. Therefore, developing a novel asymmetric duplex encoder, with ViT and CNN employed separately in each branch, for RGB-T scene parsing, stands as an underexplored area of research deserving greater attention.
@@ II-A Single-Modal Scene Parsing
The advent of the fully convolutional network (FCN) [26] catalyzed a wave of research on CNN-based scene parsing networks, which typically utilize convolutions for pixel-wise classification and explore various modules and architectures to enhance overall performance. For instance, the DeepLab series [27, 28] leverages atrous spatial pyramid pooling module to capture multi-scale context information, while the feature pyramid network (FPN) [29] effectively fuses multi-scale features in a fully convolutional fashion for robust dense prediction. Nonetheless, CNNs suffer from inherent limitations, e.g., constrained receptive fields, which impede their effectiveness in comprehensively modeling global context.
Transformers [30], originally proposed for natural language processing, have revolutionized computer vision due to their capability to model long-range dependencies. As a representative example, segmentation Transformer (SETR) [31] introduces a ViT-based encoder for scene parsing, enabling long-range dependency modeling via self-attention mechanisms. Another approach, SegFormer [32], employs a hierarchical encoder to achieve multi-scale feature encoding within the Transformer framework, outperforming SETR in handling objects of varying sizes. Additionally, drawing upon earlier works [33, 34], the MaskFormer [35, 36] series formulates scene parsing as a mask classification problem. These networks utilize a Transformer decoder to refine object queries and generate pixel-level predictions by decoding these queries into a set of masks. Given the impressive performance demonstrated by the MaskFormer architecture in dense prediction tasks, we adopt a similar network design for scene parsing in this study.
@@ II-B Data-Fusion Scene Parsing
To address this drawback, recent works [21, 9, 8, 10, 6, 20, 7] have resorted to more advanced, learnable feature fusion strategies. For instance, FEANet [21] and GMNet [9] introduce attention-based modules to dynamically select important heterogeneous features. Their subsequent research endeavor, CMX [8], employs average and max pooling operations along with the attention modules to recalibrate heterogeneous features at the spatial and channel dimensions. Similarly, CAINet [10] employs different attention modules to model global context and aggregate local semantics across various spatial scales of RGB-T feature maps. Moreover, the ABMDRNet series [6, 20] introduces novel modality-wise bridging-then-fusing frameworks, capable of reconstructing pseudo images of one modality using the features extracted from the other modality. This approach successfully reduces modality discrepancy and generates more discriminative fused features. Additionally, inspired by the masked autoencoder (MAE) [38], CRM-RGBTSeg [7] randomly masks fixed-size regions within one modality while simultaneously providing complementary regions from another modality to duplex encoders, thus preventing over-reliance on either modality. In this article, we employ a novel asymmetric duplex encoder to capture both local spatial priors and global context from the RGB-T data. These heterogeneous features are progressively fused for more robust scene parsing.
@@ III-A Overall Workflow
Unlike previous works that utilize symmetric duplex encoder structures, we innovatively introduce an asymmetric, hybrid architecture (based on both a Transformer-based VFM and a lightweight CNN). It effectively leverages the complementary strengths of both modalities, thereby achieving deeper cross-modal feature fusion and further improving the overall performance. This is realized by dividing the ViT encoder into four blocks and inserting the PHFI to form four encoding stages. The RGB image $\boldsymbol{I}^{R}\in\mathbb{R}^{H\times W\times 3}$ and its corresponding thermal image $\boldsymbol{I}^{T}\in\mathbb{R}^{H\times W\times 3}$ are first fed into the CSPD, extracting the cross-modal spatial prior $\boldsymbol{F}^{P}_{1}\in\mathbb{R}^{(\sum_{i=2}^{4}{\frac{HW}{{S_{i}}^{2}})}\times D}$, which serves as one of the inputs to following stages, where $S_{i}=2^{i+1}$ ($i\in[2,4]\cap\mathbb{Z}$) denotes the corresponding stride number. Afterwards, $\boldsymbol{I}^{R}$ is tokenized to form the context feature $\boldsymbol{F}^{V}_{1}\in\mathbb{R}^{\frac{HW}{16^{2}}\times D}$, which serves as the other input to following stages. $\boldsymbol{F}^{V}_{1}$ and $\boldsymbol{F}^{P}_{1}$ are then fused in a dual-path, progressive manner using the two key components of our developed PHFI: (1) the global-local context aggregator (GLCA) and (2) the complementary context generator (CCG). Meanwhile, fused features undergo global context encoding in four ViT blocks. This design enables the RGB features to retain fine-grained local semantics while effectively capturing global context through the powerful VFM. Eventually, the fused multi-scale features $\mathcal{F}^{F}=\{\boldsymbol{F}^{F}_{\frac{1}{S_{j}}}\in\mathbb{R}^{\frac{H}{S_{j}}\times\frac{W}{S_{j}}\times D}\}$, where $S_{j}=2^{i+1}$ ($j\in[1,4]\cap\mathbb{Z}$) from the four encoding stages are obtained and then fed into the subsequent mask classification decoder to generate semantic predictions $\boldsymbol{M}^{P}\in\mathbb{R}^{H\times W}$, which are also utilized by an auxiliary local semantics enhancement task during model training.
@@ III-B3 Progressive Heterogeneous Feature Integrator
where $\mathrm{MHA}(\cdot)$ represents the multi-head attention (MHA) operation (implemented based on multi-scale deformable attention [36] to reduce computations) and $\mathrm{LN}(\cdot)$ represents the layer normalization operation. Following common practices in attention modules such as [45, chen2022vision, 43], we introduce a learnable coefficient $\kappa_{i}$ to dynamically adjust the weight of the MHA, enabling flexible fusion of the local semantics-enhanced features into the VFM. After the $i$-th ViT block, the output feature maps $\boldsymbol{F}^{V}_{i+1}$ with enriched global context are yielded. $\boldsymbol{F}^{V}_{i+1}$ is then fed into the CCG to fuse its global context with the spatial prior $\boldsymbol{F}^{P}_{i}$. In this process, $\boldsymbol{F}^{P}_{i}$ serves as the query, while $\boldsymbol{F}^{V}_{i+1}$ acts as the keys and values to perform MHA within the CCG as follows:
@@ III-E Loss Function
We use a combination of the cross-entropy (CE) loss $\mathcal{L}_{\text{cls}}$ for class prediction, along with the binary CE loss $\mathcal{L}_{\text{bce}}$ and the dice loss $\mathcal{L}_{\text{dice}}$ for mask prediction. For our local semantics enhancement task, we directly calculate the standard CE loss $\mathcal{L}_{\text{ce}}$ over its semantic predictions. Therefore, the overall loss function can be expressed as follows:
@@
Following [36], we set $\lambda_{\text{bce}}=5.0$, $\lambda_{\text{dice}}=5.0$, $\lambda_{\text{cls}}=2.0$, and $\lambda_{\text{ce}}=0.4$, respectively, and we set $\lambda_{\text{cls}}=0.1$ for the “no object” category. Additionally, we employ bipartite matching [47] to determine the least-cost assignment during model training.
@@ IV-C Implementation Details
Our model was trained for 200 epochs on an NVIDIA RTX 3090 GPU, utilizing the AdamW optimizer [60] with an initial learning rate of $2\times 10^{-4}$ and a weight decay of $5\times 10^{-2}$. Following the common practice employed in [16], we apply a layer-wise learning rate decay of $0.9$ for our VFM encoder. For the experiments conducted on the NYU dataset, we employ RGB images and the HHA encoding of depth images as model input. We conduct all ablation studies on the widely utilized MFNet dataset, with the auxiliary task omitted to ensure training efficiency. The performance improvement (an increase in mIoU by $0.6\%$) between Sect. IV-D and Sect. IV-E demonstrates the effectiveness of the auxiliary task.
@@ IV-D Comparisons with SoTA Scene Parsing Networks
Specifically, our proposed HAPNet achieves the highest mIoU across all datasets, outperforming SoTA methods by $0.1$-$21.8\%$ on the MFNet dataset, $1.0$-$32.0\%$ on the PST900 dataset, and $2.4$-$33.6\%$ on the KP day-night dataset, respectively. These results demonstrate the robustness and effectiveness of HAPNet for RGB-T scene parsing in various scenarios, ranging from complex urban driving scenes to challenging underground scenes. Notably, HAPNet outperforms CRM-RGBTSeg [7], another SoTA Transformer-based approach based on the mask classification paradigm, demonstrating the efficacy of our hybrid, asymmetric architecture in heterogeneous feature fusion. The qualitative results also demonstrate the superior capability of HAPNet in handling various challenging scenes with poor lighting conditions.